Veronica has observed 200 bicycles at an intersection and found that 80% did not make a complete stop. Veronica tells you that the margin of error was 4%. What was the confidence level of the study?
Answer by stanbon(75887)

Veronica has observed 200 bicycles at an intersection and found that 80% did not make a complete stop. Veronica tells you that the margin of error was 4%. What was the confidence level of the study?
------
ME = z*sqrt(pq/n)
----
0.04 = z*sqrt[0.8*0.2/200)
---
0.04 = z*0.0283
---
z = 1.4142
======================================
normalcdf(1.4142,100) = 0.0786
---
Confidence level = 1 - 2*0.0786 = 84.27%
============================================
Cheers,
Stan H.
=============